OpenStudy (anonymous):

Graph the expression sec^ 2 x - sec^ 2 x sin^ 2 x on your calculator. Determine what constant or single function is equivalent to the given expression.

OpenStudy (anonymous):

factor out the common thing and see if you can simplify what remains

OpenStudy (anonymous):

So it would be 1/sin ^2x -1/sin^2x sin^2x/1

OpenStudy (anonymous):

you need to use the equation editor... also, sec x is not 1/sin x

OpenStudy (anonymous):

\[\sec ^{2}x- \sec ^{2}x \sin ^{2}x\] is this what you have?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so factor first and write what you get... i'll wait and use the equation editor

OpenStudy (anonymous):

please

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos^2x } - \frac{ 1 }{ \cos^2x}\sin ^2x\]

OpenStudy (anonymous):

not what I was looking for quite yet... you should factor first. good job on using the editor!

OpenStudy (anonymous):

What did u mean by factoring ?

OpenStudy (anonymous):

do you know the distributive property? usually students only think of it as going one way, but it goes both. no pun intended. a(b+c) => ab + ac this is how most think of the distributive property. but it also works like this ab + ac => a(b+c) when it's done like this it's called factoring. look familar?

OpenStudy (anonymous):

Ok but how would I use in this problem

OpenStudy (anonymous):

what do both terms in your expression have in common? factor that out...

OpenStudy (anonymous):

Ok so all that will be left is sin^2x

OpenStudy (anonymous):

no... look \[\sec ^{2}x- \sec ^{2}x \sin ^{2}x= \sec ^{2}x \left( 1-\sin ^{2}x \right)\]

OpenStudy (anonymous):

what's 1-sin^2 (x) equal to? what's sec^2 (x) equal to? what do you get when you multiply them together?

OpenStudy (anonymous):

yo dude... did you get this one?